OpenStudy (ebonyserrano):

Ivan will rent a car for the weekend. He can choose one of two plans. The first plan has an initial fee of $49.98 and costs an additional $0.19 per mile driven. The second plan has an initial fee of $55.98 and costs an additional $0.14 per mile driven. How many miles would Ivan need to drive for the two plans to cost the same?

OpenStudy (ebonyserrano):

@Will.H

OpenStudy (thatonegirl_):

make an equation for the first and second and set them equal to each other. Any ideas on the first equation?

OpenStudy (will.h):

The 1st equation would be Y=0.19x + 49.98 The 2nd would be Y= 0.14x + 55.98 Use substitution method to solve 0.14x + 55.98 = 0.19x + 49.98 6 = 0.05x 120 = x

OpenStudy (ebonyserrano):

so it would be 120 miles? @Will.H

OpenStudy (will.h):

Yeah since x represents miles

OpenStudy (ebonyserrano):

thank you!